# A Unified Framework for Variable Selection in Model-Based Clustering with Missing Not at Random

### v2

@@ Acknowledgments
TrungTin Nguyen and Christopher Drovandi acknowledge funding from the Australian Research Council Centre of Excellence for the Mathematical Analysis of Cellular Systems (CE230100001). Christopher Drovandi was also supported by an Australian Research Council Future Fellowship (FT210100260). The authors would like to express their special thanks to the authors of [37, 39], in particular Professor Cathy Maugis-Rabusseau and Dr. Marie-Laure Martin, for providing the Arabidopsis thaliana transcriptome dataset.
@@ Appendix A More Details on Main Challenges and Contributions
    While full optimality of $c$ is not derived, we provide a condition for choosing $c$ to achieve a target false positive rate in relevant variables selection, given the single-decision error probability. This moves beyond purely heuristic choices for $c$ in [13], paving the way for more adaptive or theoretically grounded choices of $c$ in practice. This offers a more principled approach to selecting this single hyperparameter in the algorithm, enhancing the reliability of the selection procedure.
@@
Similar to the global GMM representation of $\mathbb{S}$$\mathbb{R}$} model under MAR, we can obtain such global form of ${\mathbb{S}}{\mathbb{R}}{\mathbb{U}}{\mathbb{W}}$ model under the same missingness pattern, which is illustrated by the following claim .
@@ D.1 Proof of Theorem 1
This implies the full set of model specifications and parameters for ${\mathcal{M}}_{A}$ and ${\mathcal{M}}_{B}$ are identical. Therefore, the SRUW model under the specified MNARz mechanism is identifiable. ∎
@@ D.2 Proof of Theorem 2
Note that $\tilde{s}_{M}$ will depend on the sizes of the blocks and it is finite with being dependent on $s_{M}$, $\rho$, and the structure of $\bm{\Omega}$}, $\bm{\beta}$}, and ${\bm{\Sigma}}_{k}$.
@@
Now we prove that the envelope function $F$, which is related to the upper and lower bounds aboe, is $h$-integrable. The true density $h({\bm{y}})$ corresponds to a Gaussian mixture model $f({\bm{y}};{\bm{\theta}}^{*}_{({\mathbb{S}}_{0},{\mathbb{R}}_{0})})$ with parameters in a compact set. The observed-data density for ${\bm{y}}^{o}$, obtained by marginalizing over the missing components, is given by
@@
Given that ${\bm{\Gamma}}\in{\mathcal{D}}_{\lvert{\mathbb{W}}\rvert}$ and is independent of any relevant variables, we have the principal sub-matrix ${\bm{\Gamma}}_{oo}$ is bounded by $s_{M}$
@@ D.3 Proof of Theorem 3
For two lemmas below, we will use the following assumptions:
@@
By Theorem 5, the observed-data likelihood under MNARz-SRUW equals the augmented observed-data likelihood under MAR for every $({\bm{\theta}},{\bm{\psi}})$ and each observation; hence the full-sample likelihoods coincide pointwise in $({\bm{\theta}},{\bm{\psi}})$. The following corollary shows the resulting estimator-level equivalence.
@@
again modulo label switching. ∎
@@ F.1 Metric Details
We outline some metrics that we use to evaluate the clustering accuracy and imputation error, which is summarized as below:
@@ F.2 Assessing Integrated MNARz Handling
We conduct an additional experiment to showcase the performance of handling MNAR pattern within variable selection framework of which the setups are similar to [62]. Datasets were simulated with $n=100$ observations, $K=3$ true clusters with proportions ${\bm{\pi}}=(0.5,0.25,0.25)$), and $D=6,9$ variables. True cluster memberships $\mathbf{Z}$} were drawn according to $\bm{\pi}$}. The complete data $\bm{Y}$} was generated as $Y_{nd}=\sum_{k^{\prime}=1}^{K}Z_{nk^{\prime}}\delta_{k^{\prime}d}+\epsilon_{nd}$, where $\epsilon_{nd}\sim\mathcal{N}(0,1)$, and $\bm{\delta}$} defined cluster-specific mean shifts with signal strength $\tau=2.31$ with $\delta_{11}=\delta_{14}=\delta_{22}=\delta_{25}=\delta_{33}=\delta_{36}=\tau$, others zero). For a specific MNAR scenario, the class-specific intercept component $\psi_{k}^{z}$ was 0, and variable-specific slopes $\psi_{j}^{y}$ were (1.45,0.2,3,1.45,0.2,3)(1.45,0.2,-3,1.45,0.2,-3), resulting in $P(M_{ij}=1|Y_{ij},Z_{ik}=1)=\Phi(\psi_{j}^{y}Y_{ij})$.
In Figure 4 and 5, we plot the boxplots of ARI and NMRSE over 20 replications. We observe that MNAR-based approaches attain better results over methods not designed for MNAR patterns. These methods deliver competitive performance even when the true missingness mechanism is more complex (MNARy, MNARyz); thereby, supporting the conclusion in [62]. Furthermore, our framework demonstrates even slightly higher ARI and lower NRMSE in some cases compared to standard MNARz.
Figure 4: Boxplot of the ARI obtained over 20 replications of simulated data. The theheoretical ARIs are represented by a red dashed line.
@@ F.4 Computational Times and Scalability
Using the polynomial sums Using standard polynomial identities:
@@
The combination of mathematical analysis and empirical results demonstrates that our framework not only handles high missing rates effectively but actually becomes more computationally efficient as missing data increases, making it particularly suitable for challenging real-world datasets with substantial missingness.

### v3

@@ Acknowledgments
TrungTin Nguyen and Christopher Drovandi acknowledge funding from the Australian Research Council Centre of Excellence for the Mathematical Analysis of Cellular Systems (CE230100001). Christopher Drovandi was also supported by an Australian Research Council Future Fellowship (FT210100260). This paper was also supported by AISIA Research Lab, Vietnam. The authors would like to express their special thanks to the authors of [37, 39], in particular Professor Cathy Maugis-Rabusseau and Dr. Marie-Laure Martin, for providing the Arabidopsis thaliana transcriptome dataset.
@@ Appendix A More Details on Main Challenges and Contributions
    While full optimality of $c$ is not derived, we provide a condition for choosing $c$ to achieve a target false positive rate in relevant variable selection, given the single-decision error probability. This moves beyond purely heuristic choices for $c$ in [13], paving the way for more adaptive or theoretically grounded choices of $c$ in practice. This offers a more principled approach to selecting this single hyperparameter in the algorithm, enhancing the reliability of the selection procedure.
@@
Similar to the global GMM representation of $\mathbb{S}$$\mathbb{R}$} model under MAR, we can obtain such a global form of ${\mathbb{S}}{\mathbb{R}}{\mathbb{U}}{\mathbb{W}}$ model under the same missingness pattern, which is illustrated by the following claim.
@@ D.1 Proof of Theorem 1
This implies the full set of model specifications and parameters for ${\mathcal{M}}_{A}$ and ${\mathcal{M}}_{B}$ is identical. Therefore, the SRUW model under the specified MNARz mechanism is identifiable. ∎
@@ D.2 Proof of Theorem 2
Note that $\tilde{s}_{M}$ will depend on the sizes of the blocks, and it is finite being dependent on $s_{M}$, $\rho$, and the structure of $\bm{\Omega}$}, $\bm{\beta}$}, and ${\bm{\Sigma}}_{k}$.
@@
Now we prove that the envelope function $F$, which is related to the upper and lower bounds above, is $h$-integrable. The true density $h({\bm{y}})$ corresponds to a Gaussian mixture model $f({\bm{y}};{\bm{\theta}}^{*}_{({\mathbb{S}}_{0},{\mathbb{R}}_{0})})$ with parameters in a compact set. The observed-data density for ${\bm{y}}^{o}$, obtained by marginalizing over the missing components, is given by
@@
Given that ${\bm{\Gamma}}\in{\mathcal{D}}_{\lvert{\mathbb{W}}\rvert}$ and is independent of any relevant variables, we have that the principal sub-matrix ${\bm{\Gamma}}_{oo}$ is bounded by $s_{M}$
@@ D.3 Proof of Theorem 3
For the two lemmas below, we will use the following assumptions:
@@
By Theorem 5, the observed-data likelihood under MNARz-SRUW equals the augmented observed-data likelihood under MAR for every $({\bm{\theta}},{\bm{\psi}})$ and each observation; hence, the full-sample likelihoods coincide pointwise in $({\bm{\theta}},{\bm{\psi}})$. The following corollary shows the resulting estimator-level equivalence.
@@
again, modulo label switching. ∎
@@ F.1 Metric Details
We outline some metrics that we use to evaluate the clustering accuracy and imputation error, which are summarized below:
@@ F.2 Assessing Integrated MNARz Handling
We conduct an additional experiment to showcase the performance of handling MNAR pattern within the variable selection framework of which the setups are similar to [62]. Datasets were simulated with $n=100$ observations, $K=3$ true clusters with proportions ${\bm{\pi}}=(0.5,0.25,0.25)$), and $D=6,9$ variables. True cluster memberships $\mathbf{Z}$} were drawn according to $\bm{\pi}$}. The complete data $\bm{Y}$} was generated as $Y_{nd}=\sum_{k^{\prime}=1}^{K}Z_{nk^{\prime}}\delta_{k^{\prime}d}+\epsilon_{nd}$, where $\epsilon_{nd}\sim\mathcal{N}(0,1)$, and $\bm{\delta}$} defined cluster-specific mean shifts with signal strength $\tau=2.31$ with $\delta_{11}=\delta_{14}=\delta_{22}=\delta_{25}=\delta_{33}=\delta_{36}=\tau$, others zero). For a specific MNAR scenario, the class-specific intercept component $\psi_{k}^{z}$ was 0, and variable-specific slopes $\psi_{j}^{y}$ were (1.45,0.2,3,1.45,0.2,3)(1.45,0.2,-3,1.45,0.2,-3), resulting in $P(M_{ij}=1|Y_{ij},Z_{ik}=1)=\Phi(\psi_{j}^{y}Y_{ij})$.
In Figure 4 and 5, we plot the boxplots of ARI and NRMSE over 20 replications. We observe that MNAR-based approaches attain better results over methods not designed for MNAR patterns. These methods deliver competitive performance even when the true missingness mechanism is more complex (MNARy, MNARyz); thereby, supporting the conclusion in [62]. Furthermore, our framework demonstrates even slightly higher ARI and lower NRMSE in some cases compared to standard MNARz.
Figure 4: Boxplot of the ARI obtained over 20 replications of simulated data. The theoretical ARIs are represented by a red dashed line.
@@ F.4 Computational Times and Scalability
Using the polynomial sums:
@@
The combination of mathematical analysis and empirical results demonstrates that our framework not only handles high missing rates effectively but also becomes more computationally efficient as missing data increases, making it particularly suitable for challenging real-world datasets with substantial missingness.
@@ F.6 More on Transcriptome Dataset
Figure 9: Mean expression profiles for each of the 18 clusters, displayed in separate panels. Light region indicates irrelevant P.